Please help...
A boat moves 7 kph in still water. It travels 445 km upriver and 45 km downriver in a total time of 14 hours. What is the speed of the current?
All I know is that the distances are equal.
But I don't know how to set up the equation and how to solve it.
Thank you for the help.

Answer by checkley77(12844)

d=rt or d/r=t
45/(7+x)+4/(7-x)=14
[45(7-x)+45(7+x)]/(7+x)(7-x)=14
[315-45x+315+45x]/(49-x^2)=14
630/(49-x^2)=14 now cross multiply
14(49-x^2)=630
686-14x^2=630
-14x^2=630-686
-14x^2=-56
x^2=-56/-14
x^2=4
x=sqrt4
x=2 kmph is the speed of the current.
45/(7+2)+45/(7-2)=14
45/9+45/5=14
5+9=14
14=14